An airline transports about 800 people between Chicago and New York. A round-trip ticket is $300. It is estimated that with each $5 increase in ticket prices, 10 fewer passengers will fly the airline. What ticket price would maximize the income for the airline.
Answer by stanbon(75887)

An airline transports about 800 people between Chicago and New York. A round-trip ticket is $300. It is estimated that with each $5 increase in ticket prices, 10 fewer passengers will fly the airline. What ticket price would maximize the income for the airline.
---------
Revenue = price*#of passengers
------
R(x) = (300+5x)(800-10x)
-----
R(x) = -50x^2 +1000x + 240000
---------------------
Maximum revenue occurs when x = -b/(2a) = -1000/(-100)= 10
=========================
Price = 300+10*5 = $350
=========================
Cheers,
Stan H.
==================